OpenStudy (anonymous):

Use elimination to find the solution to the system of equations. 9x+4y=97 7x+7y=35

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

k

OpenStudy (danjs):

x = 77/5 y = -52/5

OpenStudy (anonymous):

x=6,y=1 x=9,y=4 x=8,y=25/4 x=11,y=6

OpenStudy (anonymous):

Those are answer choices

OpenStudy (danjs):

um

OpenStudy (anonymous):

What?

OpenStudy (danjs):

can you check the numbers in the question again

OpenStudy (anonymous):

oh its 7x-7y=35

OpenStudy (danjs):

x=9 y=4

OpenStudy (anonymous):

-2x + 4y = -4 x - 3y = 6 The system of equations above is graphed below. Find the solution to the system.

OpenStudy (anonymous):

https://app34.studyisland.com/pics/8212Q3mod.png

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

x = -6, y = -4 x = -6, y = -3 x = -6, y = 4 x = 0, y = -6

OpenStudy (danjs):

k

OpenStudy (danjs):

A

OpenStudy (anonymous):

Use elimination to find the solution to the system of equations. https://media.studyisland.com/cgi-bin/mimetex.cgi? \begin{array}{rclC25C25}4x\%20+\%20y%20&=&%206\\2x\%20-\%202y%20&=&%208%20\end{array}

OpenStudy (anonymous):

Okay the link is 4x+y=6 2x-2y=8

OpenStudy (anonymous):

Answer Choices: x=2,y=-2 x=1,y=3 x=-3,y=18 x=9,y=-30

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

k

OpenStudy (danjs):

A

OpenStudy (anonymous):

Which of the following best describes the solution to the system of equations below? 4x + 6y = 8 12x + 18y = 24 The system of equations has exactly one solution where x=2 and y=0 The system of equations has exactly one solution where x=1 and y=7/6 The system of equations has no solution. The system of equations has infinitely many solutions.

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

k

OpenStudy (danjs):

infinite

OpenStudy (anonymous):

Use elimination to find the solution to the system of equations. 9x+5y=69 3x-6y=0 Answer Choices: x=1,y=12 x=7,y=6/5 x=6,y=3 x=14,y=7

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

k

OpenStudy (danjs):

c

OpenStudy (anonymous):

Hey dude you here

OpenStudy (danjs):

k

OpenStudy (anonymous):

Use elimination to find the solution to the system of equations. 7x+3y=7 3x+7y=3 AnswerChoices: x=4/7,y=1 x=-2,y=9/7 x=1,y=0 x=-20,y=9

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

k

OpenStudy (danjs):

c

OpenStudy (anonymous):

Use elimination to find the solution to the system of equations. -x+3y=3 5x+6y=48 Answer Choices: x=24,y=9 x=-3,y=21/2 x=3,y=2 x=6,y=3

OpenStudy (danjs):

D

OpenStudy (anonymous):

Use elimination to find the solution to the system of equations. 4x+y=12 2x-3y=6 Answer Choices: x=5,y=-8 x=3,y=0 x=0,y=12 x=21/2,y=5

OpenStudy (danjs):

B

OpenStudy (anonymous):

Ok stay here @DanJS

OpenStudy (anonymous):

Identify all correct representations of the solution to the system below y=-x-2 y=2x-5

OpenStudy (anonymous):

the intersection of the lines x = 0 and y = -2. the intersection of the lines x = -2 and y = 0. the intersection of the lines x = 1 and y = -3.

OpenStudy (anonymous):

Ok @DanJS

OpenStudy (danjs):

ok , the first graph with the point is one of the solut9ions

OpenStudy (danjs):

the second link , i cant view it

OpenStudy (danjs):

and the intersection of the lines x = 1 and y = -3. is another solution

OpenStudy (anonymous):

Which of the following best describes the solution to the system of equations below? -x + 3y = -6 2x - 3y = 6 The system of equations has infinitely many solutions. The system of equations has exactly one solution where x = 1 and y = 2. The system of equations has exactly one solution where x = 0 and y = -2. The system of equations has no solution.

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

C

OpenStudy (anonymous):

Im done.I passed tommorow can you help me with a few things?